ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 41, 42 Belinsky, Eduard, University of Zimbabwe, Zimbabwe Strong summability in the integral norm and related questions The inequality of strong summability of multiple Fourier series in integral metric is proved for the Marcinkiewicz means. It holds for a subspace of L^{1}(T^{n}) of functions of H^{1} type. Generalization and connection with other problems are considered.